how many solution does this system have
y=5x-1
10x-2y=5
a. one solution
b.no solution
c.infinite solution
d.none of these
1
Expert's answer
2013-01-16T10:12:10-0500
y=5x-1
10x-2y=5We can substitute y=5x-1 to the secondequation
10*x-2*(5x-1)=5
10x-10x+2=5
2=5
obtained equality is false so there is no solution, right answer b.